# On an unconditional $\rm GL_{3}$ analog of Selberg’s result

### v1

@@ Abstract.
Let $S_{F}(t)=\pi^{-1}\arg L(1/2+it,F)$, where $F$ is a Hecke–Maass cusp form for $\rm SL_{3}(\mathbb{Z})$ in the generic position with the spectral parameter $\nu_{F}=\big{(}\nu_{F,1},\nu_{F,2},\nu_{F,3}\big{)}$ and the Langlands parameter $\mu_{F}=\big{(}\mu_{F,1},\mu_{F,2},\mu_{F,3}\big{)}$. In this paper, we establish an unconditional asymptotic formula for the moments of $S_{F}(t)$. Previouly, such a formula was only known under the Generalized Riemann Hypothesis. The key ingredient is a weighted zero-density estimate in the spectral aspect for $L(s,F)$ which was recently proved by the authors in [18].
@@
Moments, $\rm GL_{3}$ Hecke–Maass cusp forms, Selberg’s limit theorem.
@@
Q. Sun was partially supported by the National Natural Science Foundation of China (Grant Nos. 12471005 and 12031008) and the Natural Science Foundation of Shandong Province (Grant No. ZR2023MA003).
@@ 1. Introduction
It is a well-established fact that the function $S(t)=\frac{1}{\pi}\arg\zeta\left(\frac{1}{2}+it\right)$ is associated with the number of non-trivial zeros of the Riemann zeta function $\zeta(s)$ up to the height $t>0$. The argument is defined by continuous variation along the horizontal segment from $\infty+it$ to $1/2+it$, commencing at $\infty+it$ with the value 00. The intricate behavior of $S(t)$ represents a highly intriguing topic and has been explored by numerous number theorists (see, e.g., [1, 6, 12, 22]).
@@
In 1940s, Selberg [19, 20] examined the distribution of $S(t)$ and demonstrated that for any $n\in\mathbb{N}$,
$\frac{1}{T}\int_{T}^{2T}S(t)^{2n}\mathrm{d}t=\frac{(2n)!}{n!(2\pi)^{2n}}(\log\log T)^{n}+O\big{(}(\log\log T)^{n-1/2}\big{)},$ (1.1)
@@
which implies that, in an average sense, $|S(t)|$ has an order of magnitude $\sqrt{\log\log T}$. Additionally, for the Dirichlet $L$-function $L(s,\chi)$ with primitive $\chi$, Selberg [21] studied the behavior of $S(t,\chi)=\frac{1}{\pi}\arg L\big{(}\frac{1}{2}+it,\chi\big{)}$ and obtained, for prime modulus $q$ and any fixed $t>0$,
$\frac{1}{q}\,\,\sideset{}{{}^{\star}}{\sum}_{\chi\bmod q}S(t,\chi)^{2n}=\frac{(2n)!}{n!(2\pi)^{2n}}(\log\log q)^{n}+O\big{(}(\log\log q)^{n-1/2}\big{)},$ (1.2)
@@
where $\star$ indicates that the summation runs over the primitive characters $\chi(\bmod\,q)$. It should be noted that the counterparts of (1.1) and (1.2) for the odd-order moments can also be derived by making minor adjustments to Selberg’s approach.
@@
For the $\rm GL_{2}$ family of $L$-functions, Hejhal–Luo [8] first considered the generalization of (1.2) to the case of eigenfunctions $\psi$ of the hyperbolic Laplacian $\Delta$ for the full modular group. More precisely, let $L(s,\psi)$ be the standard Hecke–Maass $L$-function associated to $\psi$ and define
$S(t,\psi)=\frac{1}{\pi}\arg L\big{(}\frac{1}{2}+it,\psi\big{)}.$
@@
They established the analogue of (1.1) when $\psi$ is an Eisenstain series. And when $\psi$ is the Hecke–Maass cusp form $u_{j}$ of $\rm SL_{2}(\mathbb{Z})$ with the first Fourier coefficient $\nu_{j}(1)$ and Laplacian eigenvalue $\lambda_{j}=\frac{1}{4}+t_{j}^{2}$ ($t_{j}\geq 0$), under Generalized Riemann Hypothesis (GRH), they showed that for fixed $t,\beta>0$ and any large parameter $T>0$,
$\begin{split}\sum_{j=1}^{\infty}e^{-\beta(t_{j}-T)^{2}}|\nu_{j}(1)|^{2}S(t,u_{j})^{2n}=\frac{2\pi^{-3/2}}{\sqrt{\beta}}\frac{(2n)!}{n!(2\pi)^{2n}}T(\log\log T)^{n}+O_{t,n,\beta}\big{(}T(\log\log T)^{n-1/2}\big{)}.\end{split}$
@@
Recently, Sun–Wang [17] reestablished this asymptotic formula by means of the zero-density estimate of weighted type instead of assuming GRH. Moreover, for the holomorphic case, Liu–Shim [14] and Sun–Wang [16] successively obtained unconditional analogues in the weight aspect and in the level aspect, respectively.
For the $\rm GL_{3}$ family of $L$-functions, Liu–Liu [13] explored the average behavior of the analogue function
$S_{F}(t):=\frac{1}{\pi}\arg L(1/2+it,F)$
@@
defined for Hecke–Maass cusp forms $F$ for $\rm SL_{3}(\mathbb{Z})$. To state the result, we need to introduce some notations. Let $F$ run over an orthogonal basis of Hecke–Maass cusp forms for $\rm SL_{3}(\mathbb{Z})$. For each $F$, we let $\mu_{F}=\big{(}\mu_{F,1},\mu_{F,2},\mu_{F,3}\big{)}$ be its Langlands parameter and have that $\mu_{F}$ is a point in the region
@@
in the Lie algebra $\mathfrak{sl}_{3}(\mathbb{C})$. Also, the cusp form $F$ has the spectral parameter $\nu_{F}=\big{(}\nu_{F,1},\nu_{F,2},\nu_{F,3}\big{)}$, where
@@
We will use the coordinates $\mu$ and $\nu$ simultaneously. Note that $\big{(}\nu_{F,1}-\frac{1}{3},\nu_{F,2}-\frac{1}{3}\big{)}$ is the type of $F$ as defined in Goldfeld’s book [7]. Assume that $F$ has the normalized Fourier coefficients $A_{F}(m,n)$ with $A_{F}(1,1)=1$. The $L$-function associated with $F$ is defined by
@@
and has the Euler product
$\begin{split}L(s,F)&=\prod_{p}\left(1-\frac{A_{F}(1,p)}{p^{s}}+\frac{A_{F}(p,1)}{p^{2s}}-\frac{1}{p^{3s}}\right)^{-1}\\ &=\prod_{p}\big{(}1-\alpha_{F,1}(p)p^{-s}\big{)}^{-1}\big{(}1-\alpha_{F,2}(p)p^{-s}\big{)}^{-1}\big{(}1-\alpha_{F,3}(p)p^{-s}\big{)}^{-1}.\end{split}$
@@
$\begin{split}\begin{cases}\alpha_{F,1}(p)\alpha_{F,2}(p)\alpha_{F,3}(p)=1\\ \alpha_{F,1}(p)+\alpha_{F,2}(p)+\alpha_{F,3}(p)=A_{F}(1,p)\\ \alpha_{F,1}(p)\alpha_{F,2}(p)+\alpha_{F,2}(p)\alpha_{F,3}(p)+\alpha_{F,3}(p)\alpha_{F,1}(p)=A_{F}(p,1),\end{cases}\end{split}$ (1.5)
and by Kim–Sarnak [10] that
@@
$\Lambda(s,F)=L(s,F)\prod_{j=1}^{3}\Gamma_{\mathbb{R}}(s-\mu_{F,j})=L(1-s,\widetilde{F})\prod_{j=1}^{3}\Gamma_{\mathbb{R}}(1-s+\mu_{F,j})=\Lambda(1-s,\widetilde{F}),$
@@
Let $\mu_{0}=(\mu_{0,1},\mu_{0,2},\mu_{0,3})$ and $\nu_{0}=(\nu_{0,1},\nu_{0,2},\nu_{0,3})$, and satisfy the corresponding relations (1.3) and (1.4). We also assume
@@
Throughout this paper, we let $M\!=\!T^{1-\eta}$ for some very small $20<\eta<1/$ 2. Define a test function $h_{T,M}(\mu)$ (depending on $\mu_{0}$) for $\mu=(\mu_{1},\mu_{2},\mu_{3})\in\mathbb{C}^{3}$ by
@@
where $\psi(\mu)=\exp\left(\mu_{1}^{2}+\mu_{2}^{2}+\mu_{3}^{2}\right)$ and
$P(\mu)=\prod_{0\leq n\leq A}\prod_{j=1}^{3}\frac{\left(\nu_{j}-\frac{1}{3}(1+2n)\right)\left(\nu_{j}+\frac{1}{3}(1+2n)\right)}{|\nu_{0,j}|^{2}}$
@@
$\mathcal{W}:=\left\{I,\;w_{2}=\left(\begin{smallmatrix}1&&\\ &&-1\\ &1&\end{smallmatrix}\right),\;w_{3}=\left(\begin{smallmatrix}&1&\\ -1&&\\ &&1\end{smallmatrix}\right),\;w_{4}=\left(\begin{smallmatrix}&1&\\ &&1\\ 1&&\end{smallmatrix}\right),\;w_{5}=\left(\begin{smallmatrix}&&1\\ 1&&\\ &1&\end{smallmatrix}\right),\;w_{6}=\left(\begin{smallmatrix}&&1\\ &-1&\\ 1&&\end{smallmatrix}\right)\right\}$
is the Weyl group of $\rm SL_{3}(\mathbb{R})$. The function $h_{T,M}(\mu)$ localizes at a ball of radius $M$ about $w(\mu_{0})$ for each $w\in\mathcal{W}$. Moreover, we define the spectral measure on the hyperplane $\mu_{1}+\mu_{2}+\mu_{3}=0$ by
@@
$\text{spec}(\mu):=\prod_{j=1}^{3}\left(3\nu_{j}\tan\left(\frac{3\pi}{2}\nu_{j}\right)\right)\quad\mbox{and}\quad\mathrm{d}\mu=\mathrm{d}\mu_{1}\mathrm{d}\mu_{2}=\mathrm{d}\mu_{2}\mathrm{d}\mu_{3}=\mathrm{d}\mu_{1}\mathrm{d}\mu_{3}.$
By trivial estimate, we have
@@
Fix any $t>0$ and $n\in\mathbb{N}$, under GRH, Liu and Liu [13] established that, for large positive $T$,
@@
where $L(1,\text{Ad}^{2}F)$ denotes the adjoint square $L$-function of $F$. In this paper, we are devoted to proving the asymptotic formula (1.9) without assuming the GRH. We introduce the normalizing factor
$\mathcal{N}_{F}:=\|F\|^{2}\prod_{j=1}^{3}\cos\left(\frac{3}{2}\pi\nu_{F,j}\right).$
By Rankin–Selberg theory in combination with Stade’s formula (see e.g., Blomer [3, Section 4] and Li [11, Theorem 2]), it is easy to see that
$\begin{split}\mathcal{N}_{F}\asymp\mathop{\operatorname{Res}}_{s=1}L(s,F\times\widetilde{F})\ll\|\mu_{F}\|^{\varepsilon},\end{split}$
where $L(s,F\times\widetilde{F})$ is the Rankin–Selberg square of $F$ and the implied constants depending at most on $\varepsilon$. Our main result is the following theorem.
@@
Let $t>0$ be fixed and $T$ be any large parameter, we have
@@
Moreover, we define the following probability measure $\mu_{TM}$ on $\mathbb{R}$ by
$\begin{split}\mu_{TM}(E):=\left(\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\mathbf{1}_{E}\left(\frac{S_{F}(t)}{\sqrt{\log\log T}}\right)\right)\bigg{/}\left(\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\right),\end{split}$
@@
For $F$ a Hecke–Maass cusp form of $\rm SL_{3}(\mathbb{Z})$, on an average sense, our result indicates that $|S_{F}(t)|$ has order of magnitude $\sqrt{\log\log T}$ under the assumption (1.7). In addition, as $T\rightarrow\infty$, the ratios $S_{F}(t)/\sqrt{\log\log T}$ converges in distribution to a normal distribution of mean 00 and variance $(2\pi^{2})^{-1}$.
Notation Throughout the paper, $\varepsilon$ is an arbitrarily small positive real number, which may be different at each occurrence. As usual, $e(z)=\exp(2\pi iz)=e^{2\pi iz}$ and $\delta_{m,n}$ is the Kronecker symbol that detects $m=n$. We write $f=O(g)$ or $f\ll g$ to mean $|f|\leq cg$ for some unspecified positive constant $c$. The symbol $y\asymp Y$ is to mean that $c_{1}Y\leq|y|\leq c_{2}Y$ for some positive constants $c_{1}$ and $c_{2}$.
Structure of this paper. This paper is organized as follows. In Sect. 2, we give some preliminary results needed for the proof. In Sect. 3, we establish some lemmas and provide a crucial approximation of $S_{F}(t)$. In Sect. 4, we give the detailed proof of Theorem 1.1.

### v2

# On an unconditional $\mathrm{GL}_{3}$ analog of Selberg’s result

@@ Abstract.
Let $F$ be a Hecke–Maass cusp form for $\mathrm{SL}_{3}(\mathbb{Z})$ with the Langlands parameter $\mu_{F}=\big(\mu_{F,1},\mu_{F,2},\mu_{F,3}\big)$ and the associated $L$-function $L(s,F)$. Define $S_{F}(t)=\pi^{-1}\arg L(1/2+\mathrm{i}t,F)$. When $\mu_{F}$ is in generic position, we establish an unconditional asymptotic formula for the moments of $S_{F}(t)$. Previously, such a formula was only known to hold under the Generalized Riemann Hypothesis. The key ingredient is a weighted zero-density estimate in the spectral aspect for $L(s,F)$, which has recently been proved by Sun and Wang in arXiv:2412.02416.
@@
Moments, $\mathrm{GL}_{3}$ Hecke–Maass cusp forms, Selberg’s limit theorem.
@@
†  Corresponding author
@@ 1. Introduction
Let $L(s,F)$ be an $L$-function and define
$\begin{split}S_{F}(t)=\frac{1}{\pi}\arg L\Big(\frac{1}{2}+\mathrm{i}t,F\Big).\end{split}$
Here, the argument is defined by continuous variation along the horizontal segment from $\infty+\mathrm{i}t$ to $1/2+\mathrm{i}t$, commencing at $\infty+\mathrm{i}t$ with the value 0. It is very interesting to study the function $S_{F}(t)$ in number theory. For example, in the case of the Riemann zeta function, the classical Riemann–von Mangoldt formula states that the number of zeros $\rho=\beta+\mathrm{i}\gamma$ of $\zeta(s)$ (counted with multiplicity) in the rectangle $\{(\beta,\gamma)\mid 0<\beta<1,0<\gamma\leq t\}$ is given by
$\frac{t}{2\pi}\log\frac{t}{2\pi\mathrm{e}}+\frac{7}{8}+\frac{1}{\pi}\arg\zeta(1/2+\mathrm{i}t)+O\Big(\frac{1}{t}\Big),$
which implies that the function $S_{\zeta}(t)=\pi^{-1}\arg\zeta(1/2+\mathrm{i}t)$ is closely related to the nontrivial zeros of $\zeta(s)$. As such, the fascinating behavior of $S_{\zeta}(t)$ has been intensively studied by many number theorists (see, e.g., [1, 24, 5, 8] and the references therein). On the other hand, the function $S_{F}(t)$ has a close connection with random matrix theory (see [2]). In fact, the behavior of $S_{F}(t)$ is analogous to the extreme value statistics of the characteristic polynomials of random matrices.
@@
To understand more about $S_{F}(t)$, an effective way is to calculate and analyze the various moments of $S_{F}(t)$. In the 1940s, Selberg [21, 22] examined the distribution of $S_{\zeta}(t)$ and demonstrated that for any $n\in\mathbb{N}$,
$\frac{1}{T}\int_{T}^{2T}S_{\zeta}(t)^{2n}\mathrm{d}t=\frac{(2n)!}{n!(2\pi)^{2n}}(\log\log T)^{n}+O\big((\log\log T)^{n-1/2}\big),$ (1.1)
@@
which implies that, in an average sense, $|S_{\zeta}(t)|$ has an order of magnitude $\sqrt{\log\log T}$. Additionally, for the Dirichlet $L$-function $L(s,\chi)$ with primitive $\chi$, Selberg [23] studied the behavior of $S_{\chi}(t)=\frac{1}{\pi}\arg L\big(\frac{1}{2}+\mathrm{i}t,\chi\big)$ and obtained, for prime modulus $q$ and any fixed $t>0$,
$\frac{1}{q}\,\,\sideset{}{{}^{\star}}{\sum}_{\chi\bmod q}S_{\chi}(t)^{2n}=\frac{(2n)!}{n!(2\pi)^{2n}}(\log\log q)^{n}+O\big((\log\log q)^{n-1/2}\big),$ (1.2)
@@
where $\star$ indicates that the summation runs over the primitive characters $\chi$ modulo $q$. It should be noted that the counterparts of (1.1) and (1.2) for the odd-order moments can also be derived by making minor adjustments to Selberg’s approach.
@@
For the $\mathrm{GL}_{2}$ family of $L$-functions, Hejhal and Luo [11] first considered the generalization of (1.2) to the case of eigenfunctions $\psi$ of the hyperbolic Laplacian $\Delta$ for the full modular group. More precisely, let $L(s,\psi)$ be the standard Hecke–Maass $L$-function associated to $\psi$ and define
$S_{\psi}(t)=\frac{1}{\pi}\arg L\Big(\frac{1}{2}+\mathrm{i}t,\psi\Big).$
@@
They established the analogue of (1.1) when $\psi$ is an Eisenstein series. In addition, when $\psi$ is the Hecke–Maass cusp form $u_{j}$ of $\mathrm{SL}_{2}(\mathbb{Z})$ with the first Fourier coefficient $\nu_{j}(1)$ and Laplacian eigenvalue $\lambda_{j}=1/4+t_{j}^{2}$ ($t_{j}\geq 0$), under the Generalized Riemann Hypothesis (GRH), they showed that for any fixed $t,\beta>0$ and any large positive parameter $T$,
$\begin{split}\sum_{j=1}^{\infty}\mathrm{e}^{-\beta(t_{j}-T)^{2}}|\nu_{j}(1)|^{2}S_{u_{j}}(t)^{2n}=\frac{2\pi^{-3/2}}{\sqrt{\beta}}\frac{(2n)!}{n!(2\pi)^{2n}}T(\log\log T)^{n}+O_{t,n,\beta}\big(T(\log\log T)^{n-1/2}\big).\end{split}$
@@
Recently, Sun and Wang [19] reestablished this asymptotic formula by means of a weighted zero-density estimate instead of assuming the GRH. Moreover, for the holomorphic case, Liu and Shim [16] and Sun et al. [18] successively obtained unconditional analogues in the weight aspect and in the level aspect, respectively.
@@
For the $\mathrm{GL}_{3}$ family of $L$-functions, Liu and Liu [15] explored the average behavior of the analogue function $S_{F}(t)$, where $F$ is a Hecke–Maass cusp form for $\mathrm{SL}_{3}(\mathbb{Z})$. To state their results, we need to introduce some notations. For a cusp form $F$ in an orthogonal basis of Hecke–Maass cusp forms for $\mathrm{SL}_{3}(\mathbb{Z})$, we denote its Langlands parameter by $\mu_{F}=\big(\mu_{F,1},\mu_{F,2},\mu_{F,3}\big)$, where $\mu_{F}$ is a point in the region
@@
in the Lie algebra $\mathfrak{sl}_{3}(\mathbb{C})$. Denote by $\nu_{F}=\big(\nu_{F,1},\nu_{F,2},\nu_{F,3}\big)$ the spectral parameter of $F$, where
@@
We will use the coordinates $\mu$ and $\nu$ simultaneously. Note that $\big(\nu_{F,1}-1/3,\nu_{F,2}-1/3\big)$ is the type of $F$ as defined in Goldfeld’s book [10]. Assume that $F$ is normalized so that the $(1,1)$-th Fourier coefficient $A_{F}(1,1)=1$. The $L$-function associated to $F$ is defined by
@@
which has the Euler product
$\begin{split}L(s,F)&=\prod_{p}\bigg(1-\frac{A_{F}(1,p)}{p^{s}}+\frac{A_{F}(p,1)}{p^{2s}}-\frac{1}{p^{3s}}\bigg)^{-1}\\ &=\prod_{p}\big(1-\alpha_{F,1}(p)p^{-s}\big)^{-1}\big(1-\alpha_{F,2}(p)p^{-s}\big)^{-1}\big(1-\alpha_{F,3}(p)p^{-s}\big)^{-1}.\end{split}$
@@
$\begin{split}\begin{cases}\alpha_{F,1}(p)\alpha_{F,2}(p)\alpha_{F,3}(p)=1,\\ \alpha_{F,1}(p)+\alpha_{F,2}(p)+\alpha_{F,3}(p)=A_{F}(1,p),\\ \alpha_{F,1}(p)\alpha_{F,2}(p)+\alpha_{F,2}(p)\alpha_{F,3}(p)+\alpha_{F,3}(p)\alpha_{F,1}(p)=A_{F}(p,1),\end{cases}\end{split}$ (1.5)
and by Kim and Sarnak [13],
@@
[…]
@@
Let $\mu_{0}=(\mu_{0,1},\mu_{0,2},\mu_{0,3})$ and $\nu_{0}=(\nu_{0,1},\nu_{0,2},\nu_{0,3})$, which satisfy the corresponding relations (1.3) and (1.4). We also assume that $\mu_{0}$ is in generic position, i.e.,
@@
Throughout this paper, we let $M\!=\!T^{1-\eta}$ for some sufficiently small $0<\eta<1/2$. Define a test function $h_{T,M}(\mu)$ (depending on $\mu_{0}$) for $\mu=(\mu_{1},\mu_{2},\mu_{3})\in\mathbb{C}^{3}$ by
@@
where $\psi(\mu)=\exp\big(\mu_{1}^{2}+\mu_{2}^{2}+\mu_{3}^{2}\big)$ and
$P(\mu)=\prod_{0\leq n\leq A}\prod_{j=1}^{3}\frac{\big(\nu_{j}-\frac{1}{3}(1+2n)\big)\big(\nu_{j}+\frac{1}{3}(1+2n)\big)}{|\nu_{0,j}|^{2}}$
@@
$\mathcal{W}:=\Big\{I,\;w_{2}=\Big(\begin{smallmatrix}1&&\\ &&-1\\ &1&\end{smallmatrix}\Big),\;w_{3}=\Big(\begin{smallmatrix}&1&\\ -1&&\\ &&1\end{smallmatrix}\Big),\;w_{4}=\Big(\begin{smallmatrix}&1&\\ &&1\\ 1&&\end{smallmatrix}\Big),\;w_{5}=\Big(\begin{smallmatrix}&&1\\ 1&&\\ &1&\end{smallmatrix}\Big),\;w_{6}=\Big(\begin{smallmatrix}&&1\\ &-1&\\ 1&&\end{smallmatrix}\Big)\Big\}$
is the Weyl group of $\mathrm{SL}_{3}(\mathbb{R})$. The function $h_{T,M}(\mu)$ localizes at a ball of radius $M$ about $w(\mu_{0})$ for each $w\in\mathcal{W}$.
Moreover, we define the spectral measure on the hyperplane $\mu_{1}+\mu_{2}+\mu_{3}=0$ by
@@
$\text{spec}(\mu):=\prod_{j=1}^{3}\bigg(3\nu_{j}\tan\bigg(\frac{3\pi}{2}\nu_{j}\bigg)\bigg)\quad\mbox{and}\quad\mathrm{d}\mu=\mathrm{d}\mu_{1}\mathrm{d}\mu_{2}=\mathrm{d}\mu_{2}\mathrm{d}\mu_{3}=\mathrm{d}\mu_{1}\mathrm{d}\mu_{3}.$
By a trivial estimate, we have
@@
Fix any $t>0$ and $n\in\mathbb{N}$, under the GRH, Liu and Liu [15] established that, for large positive $T$,
@@
where $L(s,\text{Ad}^{2}F)$ denotes the adjoint square $L$-function of $F$. Recently, with the automorphic Plancherel density theorem of Matz and Templier, Chen et al. [9] extended their result to $\mathrm{GL}_{n}$. In this paper, we are devoted to proving the asymptotic formula (1.10) without assuming the GRH. We introduce the normalizing factor
$\mathcal{N}_{F}:=\|F\|^{2}\prod_{j=1}^{3}\cos\bigg(\frac{3}{2}\pi\nu_{F,j}\bigg).$
By Rankin–Selberg theory in combination with Stade’s formula (see e.g., Blomer [4, Section 4] and Li [14, Theorem 2]), it is easy to see that
$\begin{split}\mathcal{N}_{F}\asymp\mathrm{Res}_{s=1}L(s,F\times\widetilde{F})\ll\|\mu_{F}\|^{\varepsilon},\end{split}$
[…]
$\begin{split}L(s,F\times\widetilde{F})&=\zeta(3s)\sum_{m,n\geq 1}\frac{|A_{F}(m,n)|^{2}}{(m^{2}n)^{s}}\\ &=\prod_{p}\prod_{i=1}^{3}\prod_{j=1}^{3}\bigg(1-\alpha_{F,i}(p)\overline{\alpha_{F,j}(p)}p^{-s}\bigg)^{-1},\quad\text{for}\,\,\,\operatorname{Re}(s)>1,\end{split}$
and the implied constants depend at most on $\varepsilon$. Our main result is the following theorem.
@@
Let $t>0$ be fixed and $T$ be any large positive parameter. We have
@@
Moreover, define the following probability measure $\mu_{TM}$ on $\mathbb{R}$ by
$\begin{split}\mu_{TM}(E):=\bigg(\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\mathbf{1}_{E}\bigg(\frac{S_{F}(t)}{\sqrt{\log\log T}}\bigg)\bigg)\bigg/\bigg(\sum_{F}\frac{h_{T,M}(\mu_{F})}{\mathcal{N}_{F}}\bigg),\end{split}$
@@
For $F$ a Hecke–Maass cusp form of $\mathrm{SL}_{3}(\mathbb{Z})$, our result indicates that, as $T\rightarrow\infty$, the ratios $S_{F}(t)/\sqrt{\log\log T}$ converge in distribution to a normal distribution of mean 0 and variance $(2\pi^{2})^{-1}$.
Notation Throughout the paper, $\varepsilon$ is an arbitrarily small positive real number, which may be different at each occurrence. As usual, $e(z)=\exp(2\pi\mathrm{i}z)=\mathrm{e}^{2\pi\mathrm{i}z}$ and $\delta_{m,n}$ is the Kronecker symbol that detects $m=n$. We write $f=O(g)$ or $f\ll g$ to mean $|f|\leq cg$ for some unspecified positive constant $c$. The symbol $y\asymp Y$ is to mean that $c_{1}Y\leq|y|\leq c_{2}Y$ for some positive constants $c_{1}$ and $c_{2}$.
Structure of this paper. The rest of this paper is organized as follows. In Section 2, we give some preliminary results needed for the proof. In Section 3, we establish some lemmas and provide a crucial approximation of $S_{F}(t)$. In Section 4, we give the detailed proof of Theorem 1.1. In Section 5, we summarize our main results, highlight the significance of our work, and discuss prospects for future research.